You drop a penny from a 70-foot high bridge. How long does it take for the penny to hit the water below? Neglect the effect of air resistance, round to the nearest 1/10 second and express your answer like: 7.5 s.

Can you please help me? Thank you!

Answer by scott8148(6628)

gravitational acceleration is 32 ft/s^2

from a standing start (initial velocity is zero) ___ d = (1/2)a t^2

70 = (1/2)32 t^2

140/32 = t^2

sqrt(140/32) = t